# Exploring Tensor Network Algorithms as a Quantum-Inspired Method for Quantum Reservoir Computing

### v1

@@ II.3 Data Preprocessing, QRC Simulation, and Classical Machine Learning Methods
The embeddings were generated by computing the expectation values of the local observables at each time step. The parameters of the Rydberg Hamiltonian include a chain geometry with a scale of $11\,\mathrm{\mu m}$, the Rabi frequency $\Omega$ set to $2\pi$. The simulation was run with a time step of $0.5\,\mathrm{\mu s}$ up to a total time of $4\,\mathrm{\mu s}$.
@@
Once the quantum embeddings were obtained, we performed classification using a simple linear fitting model with the quantum embeddings as input. In addition, we also compared the performance of our quantum-inspired QRC model with a linear and nonlinear ML models using the original PCA-reduced data as input. All of our ML classification tasks were performed using the keras package [41]. The models and their parameters are described below:
  • Non-linear fitting of the original data: A 4-layer feedforward neural network (NN) model with two hidden layers and L1 regularization was trained on the PCA-reduced data. The architecture includes two dense layers with 100 units each (ReLU activation) and a dense output layer with 10 units (softmax activation). The optimizer used was Adam, and the loss function was Sparse Categorical Crossentropy.
The accuracy of the models was determined by comparing the predicted labels with the true labels of the test set, enabling a direct comparison of the performance between the quantum embeddings and the PCA embeddings.
III Results and Discussions
III.1 Scaling Analysis of Quantum Inspired QRC Method
In this section, we present a time complexity analysis to calculate the QRC embedding of a single data point in the MNIST dataset using two different methods to simulate Rydberg Hamiltonian dynamics: the TDVP-two site, and the TDVP-one site. Note that for TDVP two-site, we have chosen the maximum bond dimension to be 100.
Figure 3: Quantum inspired QRC performance for 12 qubits (12 features) with varying bond dimensions for 1D chain of Rydberg atoms. Here the quantum embeddings were generated with the two-site TDVP method at different bond lengths. This result was compared against the quantum embedding generated by the TDVP one site method. For this study, we also compared the quantum-inspired results against results from classical models (linear and NN) with classical features.
Figure 4: Model accuracy comparison between the quantum-inspired QRC methods, and classical ML algorithms. a) This graph was generated using 10000 train set and 1000 test set points. b) This graph was generated using whole dataset and k fold cross validation using $k=5$.
Figure 1 illustrates the time taken for data embedding computation across various qubit values using two different methods: TDVP-two site and TDVP-one site. The tensor network algorithm shows a steady increase in computation time as the number of qubits increases. This shows scalability better than other methods like exact diagonalization. It should be noted that the tensor network algorithm does not exhibit the same exponential increase in computation time as observed with exact diagonalization. This suggests that the tensor network approach scales more favorably with the number of qubits, making it a more viable option for practical applications involving larger quantum systems.
Among the TDVP methods, the one-site TDVP method is the most efficient and scalable method for generating data embedding with increasing qubit numbers. However, it is important to note that tensor network simulations may not accurately simulate quantum dynamics like time evolution for a high number of qubits since the bond dimensions were kept fixed during the time evolution within the one-site TDVP algorithm. Nevertheless, the embeddings generated by the one-site TDVP method still enable accurate ML performance, which will be discussed further in the next section.
III.2 Quantum Inspired QRC Model Performance
@@
To evaluate the performance of our method, we investigated the accuracy of computing the correlators using different time evolution methods, the impact of the bond dimensions on model performance, and a comprehensive comparison with various classical and quantum-inspired models using the MNIST dataset in our numerical experiments. Our findings reveal intriguing insights into the efficacy of these methods in improving machine learning tasks. To begin with, we examined the accuracy of the correlators obtained using different methods. Specifically, we considered the one-body correlator $\langle Z_{1}\rangle$ and the two-body correlator $\langle Z_{1}Z_{2}\rangle$ versus evolution time. Although the correlators derived from the tensor network algorithms may not be exact, their temporal trends closely resemble the exact results. This observation is illustrated in Fig.(2), where we compare the correlators (for an 8-qubit system) of data embeddings using the exact diagonalization method versus the TDVP one-site, and TDVP two-site methods. The TDVP two-site method, in particular, achieves a high degree of accuracy. It is important to note that for the purposes of QRC, the exact accuracy of the correlator values is not critical; the embeddings can still yield the appropriate model accuracy.
To further validate the utility of tensor network algorithms, we investigated the impact of varying bond dimensions on model accuracy using the TDVP two-site method. As shown in Fig. 3, the performance of different classical models with classical features represents the upper and lower bounds for our quantum-inspired algorithm with different bond dimensions. Our findings indicate that the bond dimension does not have a significant impact on the accuracy of the QRC. The accuracy achieved with the QRC closely matches the accuracy of the non-linear model and outperform a classical linear model demonstrating the effectiveness of the tensor network approach. Overall, this implies that quantum-inspired method like the one-site TDVP can generate adequate quantum embeddings. Despite the one-site TDVP not producing the most accurate correlators, it remains a viable option due to its efficiency and the minimal impact of bond dimension on accuracy.
Subsequently, we investigate the performance of our quantum-inspired QRC model with an increasing number of features using the one-site TDVP technique. Specifically, we compare the accuracy of three models as discussed in Section II.3: linear fitting of original data (PCA+linear), neural network model with original data (PCA+nonlinear), and linear model with quantum embeddings generated by the one-site TDVP method in the QRC algorithm. The model accuracy is illustrated in Fig.(4). It should be noted that the plot includes error bars, representing the results of experiments conducted using k-fold cross-validation with $k=5$. The analysis of Fig.(4) reveals several key insights into the performance of quantum embeddings compared to PCA embeddings. In particular, the QRC method consistently outperforms the linear model with classical features. As the number of qubits increases, the accuracy of the model also increases and then saturates. This result is expected since more features allow for a richer representation of information in the data. Importantly, the accuracy corresponding to QRC embedding using the TDVP one-site method matches that of the nonlinear model with classical features within the error range. This highlights the main advantage of the QRC approach, demonstrating that it can achieve comparable performance to more complex nonlinear models while utilizing a linear framework. The error bars are relatively small, suggesting that the results are consistent and reliable. Consequently, this demonstrates that the use of tensor network algorithms, such as the one-site TDVP method, can be useful as a quantum-inspired QRC method for datasets with a large number of features.
IV Conclusion and Outlook
This paper explored Quantum Reservoir Computing using Tensor Network algorithms to enhance the performance of classical machine learning tasks as a quantum-inspired method. By encoding data features into a Rydberg Hamiltonian and evolving the quantum system, we extracted high-quality embeddings using different TN techniques to simulate the time-evolution dynamics. In addition, this quantum-inspired algorithm was shown to be efficient enough to generate good quantum embeddings for a large number of qubits in our numerical experiments, which enhance the overall performance of classical ML models. A future extension of this quantum-inspired algorithm could consider combining TDVP techniques at one and two sites to balance precision and performance [47], increasing its applicability for different datasets. Finally, we hope that the results of this study will motivate further studies on the possible limits of using TN to study quantum dynamics and other QML applications. Specifically, in the case of QRC, since we have only explored the utility of this quantum-inspired method in the MNIST dataset, a further exploration would be to identify other datasets where the exact simulation of quantum dynamics is critical for ML performance due to the rapid increase in entanglement of quantum phases during the time evolution process [42, 43].
Acknowledgements We would like to thank Milan Kornjača, Sheng-Tao Wang, and Scott Buchholz for their feedback and discussion.
Disclaimer About Deloitte: Deloitte refers to one or more of Deloitte Touche Tohmatsu Limited (“DTTL”), its global network of member firms, and their related entities (collectively, the “Deloitte organization”). DTTL (also referred to as “Deloitte Global”) and each of its member firms and related entities are legally separate and independent entities, which cannot obligate or bind each other in respect of third parties. DTTL and each DTTL member firm and related entity is liable only for its own acts and omissions, and not those of each other. DTTL does not provide services to clients. Please see www.deloitte.com/about to learn more.
Deloitte is a leading global provider of audit and assurance, consulting, financial advisory, risk advisory, tax and related services. Our global network of member firms and related entities in more than 150 countries and territories (collectively, the “Deloitte organization”) serves four out of five Fortune Global 500®  companies. Learn how Deloitte’s approximately 460,000 people make an impact that matters at www.deloitte.com. This communication contains general information only, and none of Deloitte Touche Tohmatsu Limited (“DTTL”), its global network of member firms or their related entities (collectively, the “Deloitte organization”) is, by means of this communication, rendering professional advice or services. Before making any decision or taking any action that may affect your finances or your business, you should consult a qualified professional adviser. No representations, warranties or undertakings (express or implied) are given as to the accuracy or completeness of the information in this communication, and none of DTTL, its member firms, related entities, employees or agents shall be liable or responsible for any loss or damage whatsoever arising directly or indirectly in connection with any person relying on this communication. Copyright ©  2025 Deloitte Development LLC. All rights reserved..

### v3

@@ II.3 Data Preprocessing, QRC Simulation, and Classical Machine Learning Methods
The embeddings were generated by computing the expectation values of the local observables at each time step. Throught this paper, the simulation was run with a time step of $0.5\,\mathrm{\mu s}$ up to a total time of $4\,\mathrm{\mu s}$. A detailed analysis of the effect of total evolution time on model performance is provided in [58].
Figure 3: This plot illustrates the exponential concentration of correlators as a function of system size for varied values of $\Omega$ around the optimal region. Shown is the mean of variance of correlators across the dataset as a measure of concentration. Results are compared for both TDVP one-site and TDVP two-site methods, highlighting the stronger concentration effect observed in the two-site approach.
Once the quantum embeddings were obtained, we performed classification using a simple linear fitting model with the quantum embeddings as input. In addition, we also compared the performance of our quantum-inspired QELM model with a linear and nonlinear ML models using the original PCA-reduced data as input. All of our ML classification tasks were performed using the keras package [60]. The models and their parameters are described in [58]. The accuracy of the models was determined by comparing the predicted labels with the true labels of the test set, enabling a direct comparison of the performance between the quantum embeddings and the PCA embeddings.
In this investigation, we evaluated the performance of the quantum-inspired QELM model using various hyperparameters. For the machine learning process, we utilized 10 features and a training set comprising 10,000 data points and a test set consisting of 1,000 data points. Additionally, we employed $k$-fold cross-validation with $k=5$ to ensure robust evaluation. The standard deviation across the folds are represented as error bars in all figures.
Figure 4: Quantum inspired QELM performance for 10 qubits (10 features) with varying $\Omega$. This result was compared against the quantum embedding generated by the TDVP one site and TDVP two site method. For this study, we also compared the quantum-inspired results against results from classical models (linear and NN) with classical features.
Figure 5: This figure illustrates the time evolution of various one-body correlators for different values of $\Omega$., generated using the TDVP two site algorithm
III Results and Discussions
A central finding of our investigation is that robust and scalable quantum-inspired QELM can be realized by leveraging the interplay between disorder and entanglement in quantum dynamics, and that highly accurate quantum simulation is not necessary to achieve strong ML performance. As illustrated in Fig.(1b), tuning the Hamiltonian parameters, specifically the Rabi frequency ($\Omega$) and interatomic distance ($d$), enables direct control over the degree of disorder and entanglement in the system, which in turn governs the expressiveness of the quantum embeddings and the resulting model accuracy. In the following, we systematically analyze how disorder and entanglement affect QELM performance, and why the TDVP one-site TN method emerges as the optimal approach for scalable quantum-inspired ML. We conclude with a discussion of scalability and the impact of feature dimensionality.
III.1 Entanglement and Concentration
A central aspect of our study is understanding how entanglement, driven by the Rabi frequency $\Omega$, shapes the expressiveness of quantum embeddings and, ultimately, the performance of QELM models. We analyze this relationship through the lens of entanglement entropy, feature concentration, and model accuracy. We begin by examining the growth of entanglement entropy as a function of time and $\Omega$, as shown in Fig.(2). For small values of $\Omega$, the system remains nearly separable, and the entanglement entropy stays low throughout the evolution. As $\Omega$ increases, the entanglement entropy rises more rapidly, reflecting the enhanced quantum correlations generated by the Rydberg Hamiltonian. This trend is especially pronounced for larger $\Omega$, where the system quickly develops significant entanglement during time evolution. Thus, $\Omega$ acts as a direct control knob for the degree of entanglement in the system. The consequences of growing entanglement are further revealed by examining the phenomenon of exponential concentration, depicted in Fig.(3). Here, we plot the variance of the single-site correlators $\langle Z_{i}\rangle$ across the dataset as a function of the number of qubits for different values of $\Omega$. For the TDVP two-site method, we observe that as either the system size or $\Omega$ increases, the variance of the correlators decreases exponentially. This exponential concentration of correlators indicates that, with high entanglement, the quantum embeddings for different input records become increasingly similar effectively “collapsing” the feature space and reducing the model’s ability to distinguish between inputs. In contrast, the TDVP one-site method exhibits a much weaker concentration effect, preserving greater diversity in the embeddings even as the number of qubits grows. The impact of these effects on model performance is summarized in Fig.(4), which shows the classification accuracy of QELM models (using both TDVP one-site and two-site methods) and classical baselines as a function of $\Omega$. For small $\Omega$, the system’s dynamics are limited, and accuracy is low. As $\Omega$ increases, accuracy improves and reaches a maximum at an intermediate, optimal value of $\Omega$ the “sweet spot” where entanglement is sufficient to enrich the feature space but not so large as to induce exponential concentration. Beyond this point, further increases in $\Omega$ cause accuracy to decline, especially for the TDVP two-site method, as excessive entanglement leads to rapid feature space concentration and loss of discriminative power. The TDVP one-site method, by discounting the entanglement, avoids this detrimental effect and maintains higher accuracy at large $\Omega$.
To further elucidate the underlying dynamics, we analyze the time evolution of the single-site correlators $\langle Z_{i}\rangle$ for different $\Omega$ which is shown in Fig.(5). At small $\Omega$, the correlators exhibit minimal oscillatory behavior and remain close to their initial values, indicating limited dynamical evolution. As $\Omega$ increases, the correlators display pronounced oscillations, typically spanning the full range between 1-1 and $1$. For large $\Omega$, these oscillations are rapidly damped, and the system quickly relaxes to a steady state. This accelerated relaxation at high $\Omega$ further reduces the diversity of transient features available for learning, contributing to the observed decline in model accuracy.
Overall, our results demonstrate that moderate entanglement controlled by $\Omega$ is essential for generating expressive quantum embeddings, but excessive entanglement leads to exponential concentration and degraded model performance. The TDVP one-site method, by discounting entanglement growth, preserves feature diversity and enables robust and scalable QELM, especially in regimes where the TDVP two-site method suffers from feature space collapse.
Subsequently, to further investigate the role of entanglement on our QELM performance, we examined the effect of correlator accuracy and bond dimension in the TN simulations. While the TDVP two-site algorithm yields correlators that are significantly more accurate than those from the TDVP one-site method for more details see [58], we find that this increased accuracy does not translate into improved model performance. As illustrated in Fig.(6), varying the bond dimension in the TDVP two site method thereby increasing the precision of the quantum simulation has negligible effect on the classification accuracy of the QELM model. This observation highlights a central finding: for QELM applications, the precise accuracy of the quantum correlators is not critical. Even approximate embeddings generated by the efficient TDVP one-site method are sufficient to achieve strong model performance. This robustness further underscores the practicality and scalability of quantum-inspired approaches for ML, where computational efficiency can be prioritized without sacrificing model accuracy.
Figure 6: Quantum inspired QELM performance for 10 qubits (10 features) with varying bond dimensions with TDVP two site algorithm. Here the quantum embeddings were generated with the two-site TDVP method at different bond lengths. This result was compared against the quantum embedding generated by the TDVP one site method. For this study, we also compared the quantum-inspired results against results from classical models (linear and NN) with classical features.
Interestingly, the TDVP one-site method generates quantum states with zero entanglement, yet still achieves model accuracy comparable to the entangled states produced by the TDVP two-site method. This suggests that, beyond entanglement, another phenomenon must be responsible for the expressiveness of the quantum embeddings and the observed model performance. As we show in the following section, disorder in the quantum dynamics emerges as a key factor enabling robust and effective QELM.
Figure 7: Analysis of Disorder Effects on Model Performance These heatmaps collectively illustrate how disorder influences model performance as a function of Hamiltonian parameters: Rabi frequency ($\Omega$) and the distance between nearest-neighbor atoms. All $\Omega$ values are expressed as multiples of $2\pi$, and embeddings are generated using the TDVP one-site algorithm. Disorder is quantified using both the variance of the correlator $\langle Z_{i}\rangle$ across sites and the Edwards-Anderson order parameter. (a) Machine learning model accuracy. (b) Variance of $\langle Z_{i}\rangle$ across sites, summed over all time steps. (c) Edwards-Anderson order parameter, summed over all time steps.
III.2 Disorder and Model Expressiveness
Disorder in the quantum system, controlled by the Hamiltonian parameters, emerges as a crucial factor in determining the expressiveness of quantum embeddings and the resulting ML performance. In our QELM framework, disorder arises from the interplay between the Rabi frequency $\Omega$; the interatomic interaction strength $V_{jk}$; which is set by the distance between atoms, and the site-dependent detuning $\Delta_{i}$; which is fixed by the scaled input features. To systematically investigate the impact of disorder, we varied both $\Omega$ and the interatomic distance while keeping $\Delta_{i}$ determined by the input data. The resulting QELM model accuracy is shown as a heatmap in Fig.(7a). Notably, the model accuracy exhibits a non-monotonic dependence on both parameters: accuracy initially increases with $\Omega$ and distance, reaches an optimal value and then declines as either parameter becomes too large. This demonstrates that there exists an optimal regime in parameter space where the QELM model achieves its highest accuracy. Specifically, peak performance is observed at intermediate values of both $\Omega$ and the interatomic distance, while values that are too low or too high lead to diminished accuracy. One of the central findings of our study is that the degree of disorder in the quantum state, as tuned by these Hamiltonian parameters, has a direct and significant impact on ML performance. Increased disorder, resulting from the competition between the Rabi drive, pairwise interactions, and inhomogeneous detuning, leads to richer quantum dynamics and more diverse quantum embeddings. This, in turn, enhances the model’s ability to capture and distinguish intricate details in the input data. To quantitatively assess disorder, we compute the variance of the single-body correlators $\langle Z_{i}\rangle$ across all sites and summed over all time steps:
$\mathrm{Variance}=\sum_{t}\frac{1}{N}\sum_{i=1}^{N}\left(\langle Z_{i}(t)\rangle-\overline{\langle Z(t)\rangle}\right)^{2},$ (3)
where $\overline{\langle Z(t)\rangle}$ is the average over all sites at time $t$. Fig.(7b) shows the heatmap of this variance as a function of $\Omega$ and interatomic distance. Importantly, the regions of highest variance coincide with the regions of optimal model performance in Fig.(7a). This strong correlation supports our conclusion that maximizing disorder within an appropriate range yields more expressive quantum embeddings and thus enhances learning outcomes. This analysis was performed using the TDVP one-site method. A similar analysis for the TDVP two-site method is shown in Fig.(1b), where the highlighted regions indicate the optimal regime for model performance.
To further substantiate the relationship between disorder and model performance, we also compute the Edwards-Anderson (EA) order parameter, a well-established metric for quantifying the degree of order in disordered systems such as spin glasses [61]. In our context, a lower EA parameter indicates higher dynamical disorder in the quantum state. The EA order parameter, summed over time, is defined as:
$q=\sum_{t}\sum_{i=1}^{N}\frac{1}{N}\langle Z_{i}(t)\rangle^{2}.$ (4)
Fig.(7c) presents a heatmap of the EA parameter as a function of $\Omega$ and interatomic distance. The region where the EA parameter reaches its minimum aligns with the region of maximal model accuracy and highest variance in single-body correlators, reinforcing our conclusion that increased disorder in the quantum state leads to improved ML performance. The underlying physical picture is that all three terms in the Hamiltonian, the Rabi drive $\Omega$, pairwise interaction $V_{jk}$, and inhomogeneous detuning must be in competition to generate rich, nontrivial quantum dynamics. For intermediate values of $\Omega$, when the distance between atoms is small, the interaction term $V_{jk}$ dominates, suppressing disorder and leading to lower accuracy. As the distance increases, the strength of $V_{jk}$ decreases, allowing the disorder term to become more significant and resulting in increased accuracy. However, with further increases in distance, the pairwise interaction becomes negligible compared to the other terms, and accuracy decreases once again. Another important factor is that, for disorder to have a significant effect, all atoms must be free to transition between the ground and excited states, rather than being blockaded. This explains why the optimal distance is found to be greater than the blockade radius. The blockade radius $R_{b}$ is defined as the distance at which the Rydberg interaction energy equals the Rabi frequency and is given by $R_{b}=\left(\frac{C}{\Omega}\right)^{1/6}$, where $C$ is the interaction coefficient. For the case where $\Omega=2\pi$, $R_{b}$ is calculated to be approximately $9.75~\mu\text{m}$, while optimal performance is observed at slightly larger distances. In summary, our results demonstrate that disorder, tuned by the interplay of Hamiltonian parameters, plays a pivotal role in enabling expressive and robust quantum embeddings. This disorder driven diversity is essential for achieving high model accuracy in quantum inspired ELMs, even in the absence of significant entanglement.
III.3 Scaling Analysis
We begin our scaling analysis by examining the computational time complexity associated with generating quantum embeddings for single data points in the MNIST dataset. Specifically, we compare two TN simulation methods for evolving the Rydberg Hamiltonian: the TDVP two site method, with a maximum bond dimension set to 100, and the TDVP one site method. For this analysis, the interatomic distance is fixed at $d=11~\mu\mathrm{m}$ and the Rabi frequency at $\Omega=2\pi$. The computation time required for embedding generation as a function of the number of qubits is shown in Fig.(8).
@@ III.3 Scaling Analysis
Both TDVP methods exhibit a steady and manageable increase in computation time as the number of qubits grows, in stark contrast to the exponential scaling observed with exact diagonalization. Among the two, the TDVP one site method is notably more efficient, making it particularly well suited for practical applications involving larger quantum systems. This favorable scaling highlights the advantage of TN approaches for simulating quantum dynamics and generating quantum embeddings in ML contexts.
Figure 9: Model accuracy comparison between the quantum-inspired QRC methods, and classical ML algorithms. This graph was generated using whole dataset and k fold cross validation using $k=5$.
Having established the computational efficiency of the TN methods, we next assess the impact of increasing the number of features on model accuracy. Using the TDVP one-site method, we compare the performance of the quantum-inspired QELM model to two classical baselines, a linear model trained on PCA-reduced data (PCA+linear) and an NN trained on PCA-reduced data (PCA+nonlinear). The results, presented in Fig.(9), demonstrate that the QELM method consistently outperforms the linear model with classical features. As the number of qubits increases, the accuracy of the QELM model also increases before reaching a saturation point, reflecting the enhanced representational power provided by richer quantum embeddings. Importantly, the accuracy achieved by the QELM embedding using the TDVP one site method matches that of the nonlinear NN model with classical features, within the margin of error. This suggests that the QELM approach introduces sufficient nonlinearity through quantum dynamics, enabling a simple linear model to achieve high performance comparable to more complex classical NN. The error bars in Fig.(9), obtained from $k$-fold cross-validation with $k=5$, are relatively small, indicating the consistency and reliability of these results. It is important to note that while the TDVP one site method offers superior efficiency and scalability, TN simulations may not fully capture the true quantum dynamics for very large systems. Nonetheless, our findings show that the embeddings generated by the TDVP one site method are sufficient to enable robust and optimal ML performance.
IV Conclusion and Outlook
In this work, we investigated QELM through the lens of tensor network algorithms, presenting a quantum-inspired approach to enhance classical machine learning tasks. By encoding data features as site-dependent detunings within a Rydberg Hamiltonian and simulating the system’s time evolution, we leveraged tensor network techniques to efficiently extract high-quality quantum embeddings. Our results demonstrate that this approach enables the simulation of quantum systems with a large number of qubits on classical hardware, providing expressive embeddings. A notable advantage of the tensor network-based method is its ability to control and manage entanglement, thereby mitigating the exponential concentration problem that can hinder learning in QELM. Our numerical experiments on the MNIST dataset highlight that optimal model performance is achieved by tuning Hamiltonian parameters to maximize disorder while maintaining moderate entanglement in the quantum states.
Looking ahead, it will be important to explore how this algorithm can be applied to more complex or real-world datasets, assessing its scalability and effectiveness in practical scenarios. Additionally, further research is needed to delineate the boundaries of tensor network methods in simulating quantum dynamics for various QML applications. In particular, investigating other datasets and scenarios where the exact simulation of quantum dynamics and the associated rapid growth of entanglement plays a critical role in machine learning performance would be valuable [62, 63]. We hope these findings encourage continued exploration of tensor network approaches for advancing both quantum-inspired and quantum machine learning.
Acknowledgements We would like to thank Milan Kornjača, Sheng-Tao Wang, Mekena Mcgrew, and Scott Buchholz for their feedback and discussion. We are also grateful to Audrey Adib-Yazdi for her assistance in creating the graphics that accompany this article.
Disclaimer As used in this publication, “Deloitte” means Deloitte Consulting LLP, a subsidiary of Deloitte LLP. Please see www.deloitte.com/us/about for a detailed description of our legal structure. Certain services may not be available to attest clients under the rules and regulations of public accounting. This publication contains general information only and Deloitte is not, by means of this publication, rendering accounting, business, financial, investment, legal, tax, or other professional advice or services. This publication is not a substitute for such professional advice or services, nor should it be used as a basis for any decision or action that may affect your business. Before making any decision or taking any action that may affect your business, you should consult a qualified professional advisor.
Deloitte shall not be responsible for any loss sustained by any person who relies on this publication.
Copyright ©  2025 Deloitte Development LLC. All rights reserved..
Supplementary Material for “Harnessing Physics-Inspired Dynamics for Robust and Scalable Quantum Extreme Learning Machines”
I Details of parameters for machine learning models
  • Non-linear fitting of the original data: A 4-layer feedforward neural network (NN) model with two hidden layers and L1 regularization was trained on the PCA-reduced data. The architecture includes two dense layers with 100 units each (Rectified Linear Unit (ReLU) activation) and a dense output layer with 10 units (softmax activation). The optimizer used was Adam, and the loss function was Sparse Categorical Crossentropy.
II Effect of $V_{jk}$ Threshold
In section II.2, the interaction term $V_{jk}$ was discussed, which becomes increasingly negligible for the farthest neighbors. These terms can be disregarded if $V_{jk}$ is below a certain threshold value. In this section, we will examine how this threshold impacts model accuracy. To this end, we have plotted a graph of model accuracy across various models against the $V_{jk}$ threshold, as shown in Fig.(1).
Figure 1: Quantum inspired QELM performance for 10 qubits (10 features) with varying $V_{jk}$ threshold for Rydberg Hamiltonian. This result was compared against the quantum embedding generated by the TDVP one site and TDVP two site method. For this study, we also compared the quantum-inspired results against results from classical models (linear and NN) with classical features.
By increasing the threshold, more neighbors are incorporated, as well as additional features in the embedding. Consequently, the accuracy of the model increases as the threshold decreases. It is important to note that the time required for performing the time evolution also increases with the inclusion of added long-range interactions. To balance overall performance, we have chosen the $V_{jk}$ threshold to be $10^{-4}$ throughout the investigations.
Figure 2: Quantum-inspired QELM performance for 10 qubits (10 features) with varying total time for time evolution. This result was compared against the quantum embedding generated by the TDVP one site and TDVP two site method. For this study, we also compared the quantum-inspired results against results from classical models (linear and NN) with classical features.
@@ II Effect of $V_{jk}$ Threshold
III Total Time for Time Evolution
[…]
IV Analysis of accuracy of correlators
@@
To begin with, we examined the accuracy of the correlators obtained using different methods. Specifically, we considered the one-body correlator $\langle Z_{1}\rangle$ and the two-body correlator $\langle Z_{1}Z_{2}\rangle$ versus evolution time. Although the correlators derived from the tensor network algorithms may not be exact, their temporal trends closely resemble the exact results. This observation is illustrated in Fig.(3), where we compare the correlators (for an 8-qubit system) of data embeddings using the exact diagonalization method versus the TDVP one-site, and TDVP two-site methods. The TDVP two-site method, in particular, achieves a high degree of accuracy. It is important to note that for the purposes of QELM, the exact accuracy of the correlator values is not critical; the embeddings can still yield the appropriate model accuracy.